A 0.160-kg hockey puck is moving on an icy, frictionless, horizontal surface. At t=0 the puck is moving to the right at 3.02 m/s .

Part A)

Calculate the magnitude of the velocity of the puck after a force of 25.4 N directed to the right has been applied for 6.0×10−2 s .

V= m/s
Answer #1

given
m = 0.16 kg
vi = 3.02 m/s
A) F = 25.4 N
delta_t = 6*10^-2 s
vf = ?

use Impulse-momentum theorem

Impulse = change in momentum

F*delta_t = m*(vf - vi)

vf - vi = F*delta_t/m

vf = vi + F*delta_t/m

= 3.02 + 25.4*6*10^-2/0.16

= 12.5 m/s <<<<<<<<-----------------------Answer